OpenStudy (anonymous):

How do I solve this rational expression? (2/x) + (1/x+4)

OpenStudy (agent0smith):

\[\Large \frac {2}{x}-\frac{1}{x+4}\] You have to get a common denominator first... which we can do like this \[\Large \frac {2}{x}*\frac{ (x +4)}{ (x+4) }-\frac{1}{(x+4)}*\frac{ x }{ x }\]

OpenStudy (isaiah.feynman):

@agent0smith Actually its the addition of the fractions.

OpenStudy (anonymous):

I worked out the problem. Thanks for the help.

OpenStudy (isaiah.feynman):

Good for you!

OpenStudy (isaiah.feynman):

Now you are above addition of rational expressions.